Given the circle x^2+y^2=25 and the point A(3.-4) on the circle. Find the equation of the tangent line to the circle.
Answer by Alan3354(69443)

Given the circle x^2+y^2=25 and the point A(3,-4) on the circle. Find the equation of the tangent line to the circle.
----------
The slope at any point is -x/y --> m = 3/4
y+4 = (3/4)*(x-3)